OpenStudy (anonymous):

How do I find the slope and y intercept of the line 3x=4y+4

OpenStudy (anonymous):

Solve in terms of y The equation will be of the form y = mx +b where m is the slope & b = y-int Solving for y: 3x - 4 = 4y y = 3/4x-1 m = 3/4= slope b = -1 = y-int

OpenStudy (anonymous):

rearrange it into y=mx+b form. 3x=4y+4 4y=3x-4 y=3/4x-1 3/4 is your slope, the y-intercept is -1

OpenStudy (anonymous):

Hi, mom!! :D <3

OpenStudy (anonymous):

These rational expressions are kicking my retrice :(